whats the way or role or formula can i usr to find this question<br />

Find the mean, variance and standard deviation of the sampling distribution of the sample proportion, p-hat. <br />
<br />
p = 0.5, n = 250 <br />
<br />
p = 0.1, n = 100 <br />
1
Expert's answer
2011-07-19T05:12:39-0400
For sampling distribution of the sample proportion, p-hat with parameters p and nmean, variance and standard deviation are calculated by the following formulae:
mean = p
variance = p(1-p)/n
standard deviation = SQUARE_ROOT(variance)
Hence if

1) p = 0.5, n = 250, then

mean = p = 0.5
variance = p(1-p)/n = 0.5 * (1-0.5) / 250 = 0.25/250 = 0.0001
standard deviation = SQUARE_ROOT(variance) = SQUARE_ROOT(0.0001) = 0.01


2) p = 0.1, n = 100, then

mean = p = 0.1
variance = p(1-p)/n = 0.1 * (1-0.1) / 100 = 0.09/100 = 0.0009
standard deviation = SQUARE_ROOT(variance) = SQUARE_ROOT(0.0009) = 0.03